OpenStudy (anonymous):

what is n and x if √(x/10)=√(2n-6)?

OpenStudy (2bornot2b):

Is it that you don't like these problems, or you are just having some real problems with these sums?

OpenStudy (anonymous):

IDK how to solve it at all. Coz, its d topic 4 2nd year. Am only 1st year student.

OpenStudy (anonymous):

So, i tried to ask my other schoolmates, then they give me this site.

OpenStudy (anonymous):

so you are challenging yourself?

OpenStudy (2bornot2b):

Ok I am giving you a general idea about these problems. The trouble with these problems is with the square root. You can't solve an equation which has a square root in it. So your first intention would be to remove the square root. And that is why in most of the problems you square both sides. Once the square root is removed, proceed as you would have done for any other general equation. Try this problem, and let us know if you get stuck somewhere

jhonyy9 (jhonyy9):

sqrt(x/10)=sqrt(2n-6) squared both sides x/10=2n-6 x=20n-60 or n=(x+60)/20